# Is Channel Independent strategy optimal for Time Series Forecasting?

### 18 Dec 2023 v3

@@ 3.2 Generalization Bound of CSC
where $T^{(i)}_{SL}$, $T^{(i)}_{SU}$, $T^{(i)}_{NU}$ and $T^{(i)}_{NM}$ denote the learnable signal component, the unlearnable signal component, the memorizable noise component and the unmemorizable noise component respectively. Since we only discuss under the condition of not changing the model capacity, we can denote $T^{(i)}_{SU}+T^{(i)}_{NU}+T^{(i)}_{NM}$ as broadly defined noise $N^{(i)}$ and $T^{(i)}_{SL}$ as signal $S^{(i)}$. We denote the input matrix and the target matrix as $X^{(i)}=X^{(i)}_{S}+X^{(i)}_{N}$ and $Y^{(i)}=Y^{(i)}_{S}+Y^{(i)}_{N}$. We define a well-trained model $\mathcal{H}^{*}$ trained on training set $T^{(i)}$ as a signal error minimizer such that *$\mathcal{H}^{*}\left(X^{(i)}_{S}\right)=Y^{(i)}_{S}$. Note that $\mathcal{H}^{*}$ is not necessarily the generalization error minimizer.
@@
Theorem 3.1
Let =$\mathcal{H}=\{x\rightarrow\left\langle w,x\right\rangle|w\in\mathbb{R}^{d},\left\|w\right\|_{2}\leq B\}$ for some constant $B\geq 0$, Moreover, assume each element of $x$ is a random variable independently drawn from a zero-mean distribution with variance $\sigma^{2}$. Then, the upper bound of the Rademacher complexity of $\mathcal{H}$ is
$R_{n}\left(\mathcal{H}\right)$ $\leq B\sigma\sqrt{\frac{d}{n}}$
@@
Let $\mathcal{H}^{*}_{CSC}$ be a well-trained linear-CSC model with one layer dominating $K$ channels and $\mathcal{H}^{*}_{CI}$ a well-trained linear-CI model trained on the same $K$ channels. Assume $N^{(i)}$ is a zero-mean random variable with variance $\sigma^{2}_{i}$. With probability at least $1-\delta$,
$\mathcal{L}\left(\mathcal{H}^{*}_{CSC}\right)-\hat{\mathcal{L}}\left(\mathcal{H}^{*}_{CSC}\right)$ $\leq\mathcal{O}\left(\frac{\sqrt{\sum^{K}_{i=1}{\sigma^{2}_{i}}}}{K\sqrt{n}}\right)+\mathcal{O}\left(\sqrt{\frac{\ln{\left(1-\frac{1}{\delta}\right)}}{Kn}}\right)$
$\mathcal{L}\left(\mathcal{H}^{*}_{CI}\right)-\hat{\mathcal{L}}\left(\mathcal{H}^{*}_{CI}\right)$ $\leq\mathcal{O}\left(\frac{\sum^{K}_{i=1}{\sigma_{i}}}{K\sqrt{n}}\right)+\mathcal{O}\left(\sqrt{\frac{\ln{\left(1-\frac{1}{\delta}\right)}}{n}}\right)$
Corollary 3.2 demonstrates that the generalization bound of linear-CSC models is smaller than linear-CI models.
@@ 4.3 Whether it is best to forecast future values only using the past values of the same channel?
Figure 2: We visualize the normalized loss matrices on different datasets. upper left: the training set of weather dataset; upper right : the test set of weather dataset; lower left: the random 21 channels of the training set of traffic dataset; lower right: the random 21 channels of the test set of traffic dataset. The darkest cell in each column indicates the input channel with the lowest loss.

### 19 Dec 2023 v4

@@ 3.2 Generalization Bound of CSC
where $T^{(i)}_{SL}$, $T^{(i)}_{SU}$, $T^{(i)}_{NU}$ and $T^{(i)}_{NM}$ denote the learnable signal component, the unlearnable signal component, the memorizable noise component and the unmemorizable noise component respectively. Since we only discuss under the condition of not changing the model capacity, we can denote $T^{(i)}_{SU}+T^{(i)}_{NU}+T^{(i)}_{NM}$ as broadly defined noise $N^{(i)}$ and $T^{(i)}_{SL}$ as signal $S^{(i)}$. Then, we can denote the input matrix and the target matrix as $X^{(i)}=X^{(i)}_{S}+X^{(i)}_{N}$ and $Y^{(i)}=Y^{(i)}_{S}+Y^{(i)}_{N}$. We define a well-trained model $\mathcal{H}^{*}$ trained on training set $\hat{T}^{(i)}$ as a signal error minimizer such that *$\mathcal{H}^{*}\left(X^{(i)}_{S}\right)=Y^{(i)}_{S}$. Note that $\mathcal{H}^{*}$ is not necessarily the generalization error minimizer.
@@
Lemma 3.1
Let =$\mathcal{H}=\{x\rightarrow\left\langle w,x\right\rangle|w\in\mathbb{R}^{d},\left\|w\right\|_{2}\leq B\}$ for some constant $B\geq 0$, Moreover, assume each element of $x$ is a random variable independently drawn from a $\sigma$-sub-Gaussian distribution. Then, with probability at least $1-\delta$, the upper bound of the Rademacher complexity of $\mathcal{H}$ is
$R_{n}\left(\mathcal{H}\right)$ $\leq BC\sigma\sqrt{\frac{d}{n}}$
$C$ $=\sqrt{2\log{\frac{2}{\delta}}}$
Lemma 3.2
Let $X_{1},\ldots,X_{n}$ be independent sub-Gaussian random variables with variance proxies $\sigma^{2}_{1},\ldots,\sigma^{2}_{n}$. Define $Z=\left\langle\mathbf{o},[X_{1},\ldots,X_{n}]\right\rangle$ where $\mathbf{o}$ is a random one-hot vector following a uniform distribution over all one-hot vectors. Then, we have
$\mathrm{P}\left(\lvert Z-\mathbb{E}[Z]\rvert\geq t\right)$ $\leq 2\exp\left(-\frac{t^{2}}{2\max\{\sigma^{2}_{1},\ldots,\sigma^{2}_{n}\}}\right)$
@@
Let $\mathcal{H}^{*}_{CSC}$ be a well-trained linear-CSC model with one layer dominating $K$ channels and $\mathcal{H}^{*}_{CI}$ a well-trained linear-CI model trained on the same $K$ channels. Assume $N^{(i)}$ is a sub-Gaussian random variable with variance proxy $\sigma^{2}_{i}$. We have
$\mathcal{L}\left(\mathcal{H}^{*}_{CSC}\right)-\hat{\mathcal{L}}\left(\mathcal{H}^{*}_{CSC}\right)$ $\leq R_{CSC}+\epsilon_{CSC}$
$\mathcal{L}\left(\mathcal{H}^{*}_{CI}\right)-\hat{\mathcal{L}}\left(\mathcal{H}^{*}_{CI}\right)$ $\leq R_{CI}+\epsilon_{CI}$
With probability at least $1-\lambda$,
$R_{CSC}$ $=\mathcal{O}\left(\frac{\max\{\sigma_{1},\sigma_{2},\ldots,\sigma_{K}\}}{\sqrt{Kn}}\right)$
$R_{CI}$ $=\mathcal{O}\left(\frac{\sum^{K}_{i=1}{\sigma_{i}}}{K\sqrt{n}}\right)$
With probability at least $1-\delta$,
$\epsilon_{CSC}=\mathcal{O}\left(\sqrt{\frac{\ln{\left(\frac{1}{\delta}\right)}}{Kn}}\right)$
$\epsilon_{CI}=\mathcal{O}\left(\sqrt{\frac{\ln{\left(\frac{1}{\delta}\right)}}{n}}\right)$
Corollary 3.3 demonstrates that the generalization bound of linear-CSC models is more likely to be smaller than linear-CI models, particularly when $K$ is large, since the noise variance of similar channels typically do not differ significantly.
@@ 4.3 Whether it is best to forecast future values only using the past values of the same channel?
Figure 2: We visualize the normalized loss matrices on different datasets. upper left: the training set of weather dataset; upper right : the test set of weather dataset; lower left: the first 21 channels of the training set of traffic dataset; lower right: the first 21 channels of the test set of traffic dataset. The darkest cell in each column indicates the input channel with the lowest loss.